OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule if appropriate. If there is a more elementary method, consider using it. lim x→0 e^2x − e^−2x − 4x/x-sinx I've tried to do it w the rule but i keep getting it wrong, please help!!

OpenStudy (caozeyuan):

\[\lim_{x \rightarrow 0}e ^{2x}-e ^{-2x}-\frac{ 4x }{ x }-\sin(x)\]

OpenStudy (caozeyuan):

am I correct in transcribing your question?

OpenStudy (anonymous):

almost, the whole thing is being divided by x-sinx

OpenStudy (anonymous):

For starters, is using the rule appropriate?

OpenStudy (anonymous):

well i plugged 0 in and i ended up with 0/0

OpenStudy (caozeyuan):

@SithsAndGiggles I learned this rule in calc 1 so I guess yes?

OpenStudy (caozeyuan):

@gaba a joke first: does you name means the neurotransmitter gamma-amino-butyic acid?

OpenStudy (caozeyuan):

I see what you mean so the correct version would be \[\lim_{x \rightarrow 0}\frac{ e^{2x}-e ^{-2x}-4x }{ x-\sin(x) }\]

OpenStudy (anonymous):

gaba its one of my nicknames, but yeah jaja

OpenStudy (anonymous):

yeah that's it

OpenStudy (caozeyuan):

Do you now what exactly is lhospital rule and how to use it?

OpenStudy (caozeyuan):

If you don't then that's the end of game, but if you do it's a piece of cake

OpenStudy (anonymous):

sort of, but i keep getting the same thing over and over again, so im confused

OpenStudy (caozeyuan):

Ok, so you know the new limit is formed by taking the quotient of the derivatives of the numerator and the denominator, right?

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

so what are the derivatives we need?

OpenStudy (anonymous):

e^2x-e^-2x-4/x-cosx?

OpenStudy (caozeyuan):

OK, I have to say you kind of don't know your stuff

OpenStudy (caozeyuan):

\[\frac{ de ^{nx} }{ dx }= n*e^{nx}\]

OpenStudy (caozeyuan):

n is integer, for now

OpenStudy (caozeyuan):

you still there? what would be the correct answer?

OpenStudy (anonymous):

i got 2e^2x+2e^-2x-4 / 1-cosx

OpenStudy (caozeyuan):

but it is still 0 over 0 right?

OpenStudy (caozeyuan):

so you got to do one more round

OpenStudy (anonymous):

yeah, and i did but i kept getting that

OpenStudy (caozeyuan):

so 4e^2x+4e^2x/sin(x)

OpenStudy (caozeyuan):

wait.... that would be 8/0....WTF

OpenStudy (caozeyuan):

my brain is so rusty cuz I had my calc 1 3 years ago

OpenStudy (anonymous):

right now i have 16e^2x-16e^-2x / -sinx

OpenStudy (anonymous):

undefined right?

OpenStudy (caozeyuan):

how do you get your answer?! I didn't get any 16 or -sin

OpenStudy (anonymous):

What I meant by my comment was, does the limit satisfy the conditions for using L'Hopital's rule? The answer is yes, because substituting \(x=0\) directly gives you a proper indeterminate form: \[\frac{e^{2(0)}-e^{-2(0)}-4(0)}{0-\sin 0}=\frac{e^0-e^0-0}{0-0}=\frac{0}{0}\]

OpenStudy (anonymous):

i kept going, i had 2e^2x, and then 4e^2x, and then 8e^2x because i kept getting 0/0 with all of them

OpenStudy (caozeyuan):

But I stoped on the second trial cuz the upstair is 8

OpenStudy (anonymous):

\[\begin{align*}\lim_{x\to0}\frac{e^{2x}-e^{-2x}-4x}{x-\sin x}&=\lim_{x\to0}\frac{\dfrac{d}{dx}\left[e^{2x}-e^{-2x}-4x\right]}{\dfrac{d}{dx}\left[x-\sin x\right]}\\\\ &=\lim_{x\to0}\frac{2e^{2x}+2e^{-2x}-4}{1-\cos x}=\frac{2e^0+2e^0-4}{1-\cos0}=\frac{0}{0}\\\\ &=\lim_{x\to0}\frac{\dfrac{d}{dx}\left[2e^{2x}+2e^{-2x}-4\right]}{\dfrac{d}{dx}\left[1-\cos x\right]}\\\\ &=\lim_{x\to0}\frac{4e^{2x}-4e^{-2x}}{\sin x}=\frac{4e^0-4e^0}{\sin0}=\frac{0}{0}\\\\ &=\lim_{x\to0}\frac{\dfrac{d}{dx}\left[4e^{2x}-4e^{-2x}\right]}{\dfrac{d}{dx}[\sin x]}\\\\ &=\lim_{x\to0}\frac{8e^{2x}+8e^{-2x}}{-\cos x}=\frac{8e^0+8e^0}{\cos0}\neq\frac{0}{0}\end{align*}\]

OpenStudy (anonymous):

so then my answer is -16?

OpenStudy (anonymous):

Sorry, I meant, not negative.

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

yw